OpenStudy (anonymous):

What is 3.4% of 60? Round to the nearest hundredth.

OpenStudy (bakonloverk):

\[\frac{ 3.4 }{ 100 } \times \frac{ x }{ 60 } \]

OpenStudy (bakonloverk):

solve the proportion

OpenStudy (anonymous):

im not good with math i dont know the answer...

OpenStudy (anonymous):

Yes ! Whats the problem ?! Find with : 3.4.60/100=?

OpenStudy (bakonloverk):

3.4*60 and divide that answer by 100

OpenStudy (anonymous):

2.4 Thanks

OpenStudy (bakonloverk):

np

OpenStudy (anonymous):

@bakonloverk : Yeah !